Payback Period and NPV of Alternative Automobile Purchase
Wendy Li decided to purchase a new Honda Civic. Being concerned about environmental issues she is leaning toward a Honda Civic Hybrid rather than the completely gasoline-powered LX model. Nevertheless, she wants to determine if there is an economic justification for purchasing the Hybrid, which costs $5,100 more than the LX. Based on a mix of city and highway driving she predicts that the average gas mileage of each car is 40 MPG for the Hybrid and 30 MPG for the LX. Wendy also anticipates she will drive an average of 12,000 miles per year and that gasoline will cost an average of $3.30 per gallon over the next four years. She also plans to replace whichever car she purchases at the end of four years when the resale values of the Hybrid and the LX are predicted to be $12,000 and $8,500 respectively.

Required
a. Determine the payback period of the incremental investment associated with purchasing the Hybrid.
    Answer

years

b. Determine the net present value of the incremental investment associated with purchasing the Hybrid at a ten percent time value of money.
    Use a negative sign with your answer, if appropriate. Round answer to the nearest whole number.
    $Answer

c. Determine the cost of gasoline required for a payback period of three years on the incremental investment.

   Round answer to two decimal places

   $Answer
Answer #1
Payback period is the time period in which the initial investment is recovered
Difference in cash flows is as follows:
Year Cash flow Cumulative cash flows PVF PV
0 -5100 -5100 1 -5100
1 330 -4770 0.909091 300
2 330 -4440 0.826446 272.7273
3 330 -4110 0.751315 247.9339
4 3830 -280 0.683013 2615.942
NPV -1663.4
Payback period = NIL
Will never be paid back
NPV = Present value of cash inflows - present value of cash outflows
=-1663.4
Required savings each year = 5100/3 = $1700
Difference in Gallons per year = 100
Desired cost of gasoline = 1700/100 = $17 per gallon